One angle is thrice another. What are their measures, if they are supplementary, what are the angles?
Answer by stanbon(75887)

One angle is thrice another. What are their measures, if they are supplementary, what are the angles?
----
Equation:
x = 3y
x + y = 180
-----
3y + y = 180
4y = 180
y = 45
----
Solve for "x":
x = 3y = 3*45 = 135
----
The angles are 45 and 135
-----
Cheers,
Stan H.
=================